Consider a 1.245 g sample of glucose (C6H12O6). Calculate the moles of glucose in the sample.



1
Expert's answer
2021-04-27T07:48:39-0400

The number of moles can be calculated using the following equation:

n = m / Mr

where n - number of moles, m - mass, Mr - molecular weight.

From here:

n(C6H12O6) = m(C6H12O6) / Mr(C6H12O6) = 1.245 g / 180.156 g/mol = 0.007 mol


Answer: 0.007 mol